Three machines are filling water bottles. The machines can fill the daily quota of water bottles in 12 hours, 15 hours and 20 hours respectively. How long would it take to fill the daily quota of water bottles with all three machines working together.
Answer by checkley77(12844)

1/12+1/15+1/20=1/x (LCD=60)
(5+4+3)/60=1/x
12/60=1/5
1/5=1/x
x=5 hours to fill the quota whan all three machines are used.